OpenStudy (anonymous):

Use the substitution method to solve the linear system given. Express the answer as an ordered pair of the form (x,y) y = x + 2 3x + 2y = 9

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3x + 2(x+2) = 9 3x +2x + 4 = 9 5x + 4 = 9 5x = 5 x =1

OpenStudy (anonymous):

y = (1) + 2 y = 3

OpenStudy (anonymous):

so, x = 1, y =3

OpenStudy (anonymous):

@OM14forever